# Generative Evolutionary Meta-Solver (GEMS): Scalable Surrogate-Free Multi-Agent Learning

### v1

@@ 1 Introduction
Training AI agents in two-player zero-sum Markov games faces a similar challenge. Population-based methods such as Policy-Space Response Oracles (Psro) ([6]) maintain a growing set of $k$ policies and explicitly construct a $k\times k$ payoff matrix, where each entry records the expected outcome of a policy against another. A meta-solver then updates the distribution over policies, analogous to computing player rankings from the tournament table.
While conceptually straightforward, classical Psro suffers from three critical bottlenecks: ❶ Memory overhead: storing a new policy for each player leads to linear growth in storage; ❷ Computation overhead: filling in the full $k\times k$ payoff matrix quickly becomes infeasible; ❸ Scalability of new entries: adding a new policy requires training and storing another separate actor. Prior work has mitigated some of these costs through selective best-response training ([19]), Double Oracle methods ([11, 4]), meta-game improvements ([13],[12]), and knowledge transfer for new agents ([20, 8]). However, these approaches retain the core paradigm of explicit policy sets, leaving scalability fundamentally limited.
We introduce Gems, a surrogate-free framework that overcomes these limitations while preserving the game-theoretic guarantees of Psro. Analogous to running a large tournament without scheduling every match, Gems: ❶ Maintains a compact anchor set $Z_{t}$ of latent codes that implicitly represent active “players”; ❷ Treats the payoff matrix as conceptual and queries it only through unbiased Monte Carlo rollouts, observing a small subset of matches instead of all $k^{2}$; ❸ Updates the meta-strategy using multiplicative-weights discretization of replicator dynamics, akin to adjusting rankings based on sampled outcomes; ❹ Expands the population with an empirical-Bernstein UCB (EBUCB) oracle [9], selecting promising new “players” efficiently from a candidate pool; ❺ Trains a single amortized generator $G_{\theta}$ with an Advantage-based Best-Response objective and Trust-region Regularization (ABR–TR), eliminating the need for separate policies.
@@
Relative to classical Psro, Gems achieves: ❶ Memory efficiency: replaces $O(k)$ stored players with a single versatile generator; ❷ Computation efficiency: avoids quadratic payoff tables using Monte Carlo estimates, scaling per iteration with sampled matches and candidate pool size; ❸ Scalable new entries: EB–UCB identifies strong candidates, and ABR–TR integrates them into the generator without adding new actors; ❹ Theoretical guarantees: unbiased MC meta-gradients, instance-dependent regret for EB-UCB, external regret bounds for multiplicative-weights dynamics, and finite-population exploitability accounting for approximate best responses. Together, these advances transform the exhaustive “tournament bookkeeping” of Psro into a leaner, scalable framework, closer to how real tournaments operate, where only select matches determine rankings, and strategies adapt efficiently from sparse outcome.
@@ 2 Related Work
The seminal work of Lanctot et al. [6] introduced Psro as a general framework for multi-agent reinforcement learning. Psro iteratively expands a population of policies by training a new best-response policy against a meta-game mixture of the existing population. A core challenge of Psro is its reliance on a full $k\times k$ payoff matrix, where $k$ is the number of policies in the population. This leads to $\mathcal{O}(k^{2})$ computation overhead per iteration and $\mathcal{O}(k)$ memory overhead to store individual policies, both of which become bottlenecks in large-scale settings [18, 26]. This has motivated a line of work on more efficient variants. For example, Smith et al. [19] propose training a best-response against a single opponent, reducing computation but potentially limiting diversity. The double oracle (DO) family of algorithms, including Xdo [11] and Edo [4], focus on finding equilibria more efficiently, often by guaranteeing linear convergence or monotonic exploitability reduction [13]. However, these methods can still suffer from issues like performance oscillations or slow convergence due to the need for a full set of strategies [4]. In contrast, Gems avoids these issues entirely by replacing the explicit payoff matrix with unbiased Monte Carlo rollouts and the discrete set of policies with a single, amortized generator.
@@
Other work has addressed the computational cost of training new policies from scratch at each iteration, which is a significant bottleneck [10, 7]. To address this, Fusion-Psro [8] uses policy fusion to initialize new best responses, while Strategic Knowledge Transfer [20] explores transferring knowledge across changing opponent strategies. Our work tackles this problem differently by using a single amortized generator, which removes the need to store and train separate models for each policy.
@@
Finally, some papers have focused on improving the meta-game solution or policy diversity. Alpha-Psro [14] replaces the Nash meta-solver with $\alpha$-Rank to scale Psro to general-sum, many-player settings. A-Psro [3] introduces an advantage-based evaluation for strategy selection, providing a unified objective across zero-sum and general-sum games. Anytime Psro [13] and Self-Play Psro (SP-Psro) [12] aim to improve convergence by adding high-quality policies to the population. Other works introduce new metrics for policy diversity to ensure better approximation of the Nash Equilibrium [21, 25]. While these works aim to improve Psro, they do not fundamentally change its core structure, which still relies on an explicit population of policies and a payoff matrix.
@@
Beyond the Psro paradigm, other game-theoretic approaches exist for multi-agent learning. Counterfactual Regret Minimization (Cfr) [27] is a well-known method for extensive-form games, but it typically requires explicit state enumeration, making it challenging to scale to large or continuous domains. Some work has explored bridging the gap between Psro and Cfr, as seen in the unified perspective proposed by Wang et al. [23]. Other approaches include methods based on fictitious play, such as Fictitious Cross-Play (Fxp) [24], which combines self-play and Psro to find global Nash equilibria. However, many self-play methods lack theoretical guarantees for general-sum games and can struggle with convergence in mixed cooperative-competitive settings [24].
Our work contributes to this broader landscape by offering a surrogate-free, amortized framework that is both memory and computationally efficient. While many variants of Psro have been proposed, they have largely retained the core structure of maintaining an explicit policy population and a large payoff matrix. Gems breaks from this paradigm by using an amortized generator and unbiased Monte Carlo rollouts, thereby sidestepping the fundamental scalability issues inherent to classical Psro.
@@ 3 Proposed Method: Gems
Gems circumvents Psro ’s limitations by being surrogate-free. Instead of storing $k$ explicit actor models and computing their full payoff matrix, it maintains a single generative model $G_{\theta}$ that maps low-dimensional latent codes to policies. Consequently, the method scales to massive conceptual populations while storing only the generator and a set of latent “anchor” codes.
@@
While the core exposition focuses on two-player zero-sum games for clarity, the Gems framework extends naturally to more general settings. This extensibility is demonstrated by benchmarking Gems on two-player general-sum and, more broadly, $n$-player general-sum games. For the multi-player environments, the implementation leverages the PettingZoo library [22]. The necessary modifications to the meta-game estimators, per-player oracles, and convergence guarantees are provided in Appendix Part II.
@@ 3.1 Formal Setup and Generative Representation
    An anchor set of latent codes $Z_{t}=\{z_{1},\ldots,z_{k_{t}}\}\subset\mathbb{R}^{d}$, representing policies in a low-dimensional space.
@@
    A single generator network $G_{\theta}$ that maps a latent code $z$ to policy parameters $\varphi=G_{\theta}(z)$, thereby defining a policy $\pi_{\varphi}$.
@@
    A meta-strategy $\sigma_{t}\in\Delta_{k_{t}-1}$, which is a probability distribution over the current anchor set $Z_{t}$, representing the Nash equilibrium of the restricted game $A_{t}=\{\pi_{G_{\theta}(z)}:z\in Z_{t}\}$.
@@
$\text{Exploit}(\sigma_{t})=\max_{i\leq k_{t}}e_{i}^{\top}M\sigma_{t}\;-\;\sigma_{t}^{\top}M\sigma_{t}.$ (1)
Gems iteratively refines $\sigma_{t}$ and expands $Z_{t}$ to drive this exploitability toward zero.
@@ 3.2 Estimating the Meta-Game without the Matrix
For each anchor policy $i\in[k_{t}]$, the algorithm samples $n_{i}$ opponents $j_{s}\sim\sigma_{t}$ and executes $m$ game episodes per pair, yielding returns $Y_{i,s,\ell}\in[0,1]$. This process yields a simple yet powerful estimator for per-policy performance. To estimate the overall game value, the procedure additionally samples $B$ independent policy pairs $(i_{b},j_{b})\sim\sigma_{t}\times\sigma_{t}$ and averages their outcomes across $m$ episodes each. The resulting estimators are
@@ 3.3 Solving the Meta-Game via Optimistic Replicator Dynamics
Given the estimated payoffs $\hat{v}_{t}$ and game value $\hat{\overline{r}}_{t}$, Gems updates the meta-strategy from $\sigma_{t}$ to $\sigma_{t+1}$. Gems adopts the Optimistic Multiplicative Weights Update (OMWU) algorithm ([1]), an adaptive discretization of replicator dynamics. Rather than relying solely on the current payoff estimate $\hat{v}_{t}$, OMWU incorporates a predictive “hint” about the next payoff. Specifically, it forms the optimistic estimate $m_{t}\;=\;2\hat{v}_{t}-\hat{v}_{t-1},\text{with}\;\hat{v}_{0}=\mathbf{0}.$ The meta-strategy then updates according to
@@
followed by normalization to ensure $\sigma_{t+1}\in\Delta_{k_{t}-1}$. This optimistic step yields stronger theoretical guarantees: regret now scales with the cumulative variation of the payoff vectors rather than the iteration horizon $T$.
@@ 3.4 Finding New Strategies with a Bandit Oracle
Solving the restricted meta-game alone is insufficient; progress requires introducing new, challenging policies into the population. Gems casts this search as a multi-armed bandit problem ([15]). The “arms” comprise a finite pool of candidate latent codes $\Lambda_{t}$, and pulling an arm corresponds to evaluating the policy $G_{\theta}(z)$ against the current meta-strategy $\sigma_{t}$.
@@
Gems estimates this value, $\hat{\mu}_{t}(z)$, together with its empirical variance $\widehat{\mathrm{Var}}_{t}(z)$ via Monte Carlo rollouts. To balance exploration and exploitation efficiently, it employs an empirical-Bernstein Upper Confidence Bound (EB-UCB) ([9]), which leverages variance information for tighter bounds. A Jacobian penalty encourages smoothness in the generator’s latent space, aiding optimization. The score assigned to candidate $z$ is
$\mathrm{UCB}^{\mathrm{EB}}_{t}(z)\;=\;\hat{\mu}_{t}(z)+\sqrt{\frac{2\,\widehat{\mathrm{Var}}_{t}(z)\,\ln(3/\delta_{t})}{n_{z}m}}+\frac{3\ln(3/\delta_{t})}{n_{z}m-1}\;-\;\lambda_{J}\|JG_{\theta}(z)\|_{F}^{2},$
where $\delta_{t}=t^{-2}$ is a decaying confidence parameter. Gems selects $z_{t}^{\star}\;=\;\arg\max_{z\in\Lambda_{t}}\mathrm{UCB}^{\mathrm{EB}}_{t}(z),$ and augments the anchor set: $Z_{t+1}\leftarrow Z_{t}\cup\{z_{t}^{\star}\}$.

### v2

# Generative Evolutionary Meta-Solver (GEMS): Scalable Surrogate-Free Multi-Agent Reinforcement Learning

@@ 1 Introduction
Training AI agents in two-player zero-sum Markov games faces a similar challenge. Population-based methods such as Policy-Space Response Oracles (Psro) (Lanctot et al. (2017)) maintain a growing set of $k$ policies and explicitly construct a $k\times k$ payoff matrix, where each entry records the expected outcome of a policy against another. A meta-solver then updates the distribution over policies, analogous to computing player rankings from the tournament table.
@@
While conceptually straightforward, classical Psro suffers from three critical bottlenecks: ❶ Memory overhead: storing a new policy for each player leads to linear growth in storage; ❷ Computation overhead: filling in the full $k\times k$ payoff matrix quickly becomes infeasible; ❸ Scalability of new entries: adding a new policy requires training and storing another separate actor. Prior work has mitigated some of these costs through selective best-response training (Smith et al. (2021)), Double Oracle methods (McAleer et al. (2021); Huang et al. (2022)), meta-game improvements (McAleer et al. (2022b),McAleer et al. (2022a)), and knowledge transfer for new agents (Smith et al. (2023); Lian et al. (2024)). However, these approaches retain the core paradigm of explicit policy sets, leaving scalability fundamentally limited.
@@
Relative to classical Psro, Gems achieves ❶ memory efficiency: Gems replaces $O(k)$ stored players with a single versatile generator. ❷ Computation efficiency: Gems avoids Psro’s quadratic payoff tables using Monte Carlo estimates, scaling per iteration with the number of sampled matches and the candidate pool size. ❸ Scalable new entries: EB-UCB identifies strong candidates, and ABR-TR integrates them into the generator without adding new actors. ❹ Theoretical guarantees: unbiased MC meta-gradients, instance-dependent regret bounds for EB-UCB, external regret bounds for multiplicative-weights dynamics, and finite-population exploitability accounting for approximate best responses. Together, these advances transform the exhaustive “tournament book-keeping” of Psro into a leaner, scalable framework, closer to how real tournaments operate, where only select matches determine rankings and strategies adapt efficiently from sparse outcomes.
@@ 2 Related Work
The seminal work of Lanctot et al. (2017) introduced Psro as a general framework for multi-agent reinforcement learning. Psro iteratively expands a population of policies by training a new best-response policy against a meta-game mixture of the existing population. A core challenge of Psro is its reliance on a full $k\times k$ payoff matrix, where $k$ is the number of policies in the population. This leads to $\mathcal{O}(k^{2})$ computation overhead per iteration and $\mathcal{O}(k)$ memory overhead to store individual policies, both of which become bottlenecks in large-scale settings (Shao et al., 2024; Zhou et al., 2022). This has motivated a line of work on more efficient variants. For instance, Efficient Psro (E-psro) (Zhou et al., 2022) minimizes evaluation costs by formulating the solver as an unrestricted-restricted game. Similarly, Smith et al. (2021) propose training a best-response against a single opponent, reducing computation but potentially limiting diversity. The double oracle (DO) family of algorithms, including Xdo (McAleer et al., 2021) and Edo (Huang et al., 2022), focus on finding equilibria more efficiently, often by guaranteeing linear convergence or monotonic exploitability reduction (McAleer et al., 2022b). However, these methods can still suffer from issues like performance oscillations or slow convergence due to the need for a full set of strategies (Huang et al., 2022). In contrast, Gems avoids these issues entirely by replacing the explicit payoff matrix with unbiased Monte Carlo rollouts and the discrete set of policies with a single, amortized generator.
@@
Other work has addressed the computational cost of training new policies from scratch at each iteration, which is a significant bottleneck (Li et al., 2023). Notably, Pipeline Psro (P2SRO) (McAleer et al., 2020) mitigates this by parallelizing the training of best responses. To further address this, Fusion-Psro (Lian et al., 2024) uses policy fusion to initialize new best responses, while Strategic Knowledge Transfer (Smith et al., 2023) explores transferring knowledge across changing opponent strategies. Our work tackles this problem differently by using a single amortized generator, which removes the need to store and train separate models for each policy.
@@
Finally, some papers have focused on improving the meta-game solution or policy diversity. Alpha-Psro (Muller et al., 2019) replaces the Nash meta-solver with $\alpha$-Rank to scale Psro to general-sum, many-player settings. A-Psro (Hu et al., 2023) introduces an advantage-based evaluation for strategy selection, providing a unified objective across zero-sum and general-sum games. Anytime Psro (McAleer et al., 2022b) and Self-Play Psro (SP-Psro) (McAleer et al., 2022a) aim to improve convergence by adding high-quality policies to the population. Other works introduce new metrics for policy diversity to ensure better approximation of the Nash Equilibrium (Tang et al., 2025; Yao et al., 2023). While these works aim to improve Psro, they do not fundamentally change its core structure, which still relies on an explicit population of policies and a payoff matrix.
Table 1: Theoretical comparison of GEMS and PSRO-style methods. Let $N_{t}$ denote the cumulative number of policies found by iteration $t$. $C_{\text{eval}}$ is the cost of one payoff evaluation (e.g., one episode), and $k$ is the number of sampled opponents per iteration. Memory Scaling refers to the storage required for the meta-game state (payoff matrix / sampled entries). Eval. Cost refers to the complexity of evaluating a new policy update. While standard PSRO variants scale quadratically in meta-game memory as the population grows ($N_{t}\propto t$), GEMS maintains constant meta-game scaling ($\mathcal{O}(1)$) via its generative parameterization.
Method Payoff Matrix Memory Eval. Cost
(Meta-Game) (per iter.)
PSRO Required (Full) $\mathcal{O}(N_{t}^{2})$ $\mathcal{O}(N_{t}\,C_{\text{eval}})$
Alpha-PSRO Required $\mathcal{O}(N_{t}^{2})$ $\mathcal{O}(N_{t}\,C_{\text{eval}})$
APSRO Sampled $\mathcal{O}(N_{t})$$\mathcal{O}(N_{t}^{2})$ $\mathcal{O}(k\,C_{\text{eval}})$
EPSRO Sampled $\mathcal{O}(N_{t})$$\mathcal{O}(N_{t}^{2})$ $\mathcal{O}(k\,C_{\text{eval}})$
P2SRO Sampled / local $\mathcal{O}(N_{t})$$\mathcal{O}(N_{t}^{2})$ $\mathcal{O}(k\,C_{\text{eval}})$
NeuPL Sampled / online $\mathcal{O}(N_{t})$$\mathcal{O}(N_{t}^{2})$ $\mathcal{O}(k\,C_{\text{eval}})$
GEMS (Ours) Not Required $\bm{\mathcal{O}(1)}$ $\bm{\mathcal{O}(k\,C_{\text{eval}})}$
Note: GEMS utilizes a fixed-size latent anchor set $|Z|=K$ (where $K\ll t$ is a constant hyperparameter) to represent the population. Therefore, its meta-game memory complexity remains $\mathcal{O}(1)$ relative to training iterations $t$, whereas PSRO-based methods typically accumulate a growing population history ($N_{t}\propto t$). All methods additionally store policy parameters; the Memory Scaling column isolates meta-game (payoff/state) storage rather than network weights.
Neural Population Learning.
Recent work has aggressively explored unifying policy populations into single conditional networks to enable positive transfer and minimize storage. Neural Population Learning (NeuPL) (Liu et al., 2022b) represents an entire population of best-response policies using a single neural network that is explicitly conditioned on which strategy is being executed. Formally, this network defines a conditional policy $\pi(a\mid s,\sigma)$ that outputs a distribution over actions $a$ given the current state (or observation) $s$, where $\sigma$ is a learnable discrete strategy index identifying a particular population member. Addressing the limitations of discrete indexing, Simplex-NeuPL (Liu et al., 2022a) introduces a geometric formulation in which the policy conditions on continuous simplex vectors, enabling direct representation of mixed strategies within a single forward pass. Furthermore, NeuPL-JPSRO (Liu et al., 2024) extends this paradigm to general-sum games by combining conditional policies with Joint-PSRO objectives, effectively learning correlated joint policies that approximate coarser equilibrium concepts such as CCE. However, despite amortizing memory costs, these methods often retain the computational bottlenecks of empirical game theory, specifically the $\mathcal{O}(k^{2})$ cost of estimating pairwise payoffs to update the strategy distribution (Liu et al., 2022b). Distinct from the conditional architectures of the NeuPL family, Gems introduces a fully generative meta-solver that maps a latent geometry directly to policy parameters via a hypernetwork. This formulation allows Gems to bypass explicit population management entirely, utilizing unbiased Monte Carlo rollouts and latent-space evolutionary optimization (OMWU) to solve the game without constructing a full payoff matrix.
@@
Beyond the Psro paradigm, other game-theoretic approaches exist for multi-agent learning. Counterfactual Regret Minimization (Cfr) (Zinkevich et al., 2007) is a well-known method for extensive-form games, but it typically requires explicit state enumeration, making it challenging to scale to large or continuous domains. Some work has explored bridging the gap between Psro and Cfr, as seen in the unified perspective proposed by Wang et al. (2022). Other approaches include methods based on fictitious play, such as Fictitious Cross-Play (Fxp) (Xu et al., 2023), which combines self-play and Psro to find global Nash equilibria. However, many self-play methods lack theoretical guarantees for general-sum games and can struggle with convergence in mixed cooperative-competitive settings (Xu et al., 2023).
Our work contributes to this broader landscape by offering a surrogate-free, amortized framework that is both memory and computationally efficient. While many variants of Psro have been proposed, they have largely retained the core structure of maintaining an explicit policy population and a large payoff matrix. Gems breaks from this paradigm by using an amortized generator and unbiased Monte Carlo rollouts, thereby sidestepping the fundamental scalability issues inherent to classical Psro. As summarized in Table 1, this design removes the need to store an explicit meta-game payoff matrix and prevents meta-game state from growing with the population, while retaining a per-iteration evaluation profile comparable to sampled Psro variants.
@@ 3 Proposed Method: GEMS
Gems circumvents Psro’s limitations by being surrogate-free. [^1] In this context, “surrogate-free” signifies that GEMS does not maintain an explicit, discrete set of policies to approximate the game. Instead of storing $k$ explicit actor models and computing their full payoff matrix, it maintains a single generative model $G_{\theta}$ that maps low-dimensional latent codes to policies. Consequently, the method scales to massive conceptual populations while storing only the generator and a set of latent “anchor” codes.
@@
While the core exposition focuses on two–player zero-sum games for clarity, Gems framework extends naturally to more general settings. This extensibility is demonstrated by benchmarking Gems on two-player general-sum and, more broadly, $n$-player general-sum games. For the multi-player environments, the implementation leverages the PettingZoo library (Terry et al., 2021). The necessary modifications to the meta-game estimators, per-player oracles, and convergence guarantees are provided in Appendix II.
[…]
Figure 2: At each iteration $t$, Monte Carlo rollouts evaluate the current policy mixture under a fixed meta–strategy $\sigma_{t}$, producing estimated meta–values $\hat{v}_{t}$ (policy-to-mixture) and $\hat{\bar{r}}_{t}$ (mixture self-play). An optimistic meta-solver updates the mixture via OMWU using the hint $m_{t}=2\hat{v}_{t}-\hat{v}_{t-1}$. An EB-UCB oracle then selects a new latent anchor $z_{t}^{*}$ from the candidate set, which is incorporated through amortized generator training with a trust-region objective (ABR-TR). The anchor set $Z_{t}$ is expanded accordingly, the generator induces updated policies $\pi_{\varphi}$, and the iteration advances to $t+1$. Green ellipses denote temporal iteration boundaries rather than algorithmic operations.
Figure 2 provides a high-level overview of the Gems algorithmic loop. Rather than explicitly maintaining a growing population of policies and their full payoff matrix, Gems operates on a compact latent representation: a fixed-size generator coupled with an evolving set of latent anchors. At each iteration, Monte Carlo rollouts estimate the meta-game values under the current mixture, which are then used by an optimistic meta-solver to update the population distribution. An EB-UCB oracle selects a new latent anchor corresponding to an approximate best response, which is incorporated via trust-region amortized generator training. The anchor set is expanded, the generator induces updated policies, and the process advances to the next iteration. This closed-loop structure captures how Gems alternates between equilibrium refinement and controlled population expansion while avoiding explicit policy enumeration.
@@ 3.1 Formal Setup and Generative Representation
    An anchor set of latent codes $Z_{t}=\{z_{1},\ldots,z_{k_{t}}\}\subset\mathbb{R}^{d}$, where $d$ is the dimension of the latent space, representing policies in a low-dimensional space.
@@
[…]
@@
    A meta-strategy $\sigma_{t}\in\Delta_{k_{t}-1}$, which is a probability distribution over the current anchor set $Z_{t}$, representing the Nash equilibrium of the restricted game $A_{t}=\{\pi_{G_{\theta}(z)}:z\in Z_{t}\}$, where $\Delta_{k_{t}-1}$ is the $(k_{t}-1)$-dimensional probability simplex  [^3] The simplex $\Delta_{n}={x\in\mathbb{R}^{n+1}:x_{i}\geq 0,\sum_{i}x_{i}=1}$ denotes the space of probability distributions over $(n+1)$ elements..
@@
$\text{Exploit}(\sigma_{t})=\max_{i\leq k_{t}}e_{i}^{\top}M\sigma_{t}\;-\;\sigma_{t}^{\top}M\sigma_{t},$ (1)
where $e_{i}$ is the $i$-th standard basis vector. Gems iteratively refines $\sigma_{t}$ and expands $Z_{t}$ to drive this exploitability toward zero.
@@ 3.2 Estimating the Meta-Game without the Matrix
For each anchor policy $i\in[k_{t}]$, the algorithm samples $n_{i}$ opponents $j_{s}\sim\sigma_{t}$ and executes $m$ game episodes per pair, yielding returns $Y_{i,s,\ell}\in[0,1]$444ii indexes anchor policies in the current population, $s$ indexes sampled opponents drawn from the meta-strategy $\sigma_{t}$, and $\ell$ indexes independent Monte Carlo rollouts (episodes) used to estimate their expected returns. . This process yields a simple yet powerful estimator for per-policy performance. To estimate the overall game value, the procedure additionally samples $B$ independent policy pairs $(i_{b},j_{b})\sim\sigma_{t}\times\sigma_{t}$ and averages their outcomes across $m$ episodes each. The resulting estimators are
@@ 3.3 Solving the Meta-Game via Optimistic Replicator Dynamics
Given the estimated payoffs $\hat{v}_{t}$ and game value $\hat{\overline{r}}_{t}$, Gems updates the meta-strategy from $\sigma_{t}$ to $\sigma_{t+1}$. Gems adopts the Optimistic Multiplicative Weights Update (OMWU) algorithm (Daskalakis & Panageas (2018)), an adaptive discretization of replicator dynamics. Rather than relying solely on the current payoff estimate $\hat{v}_{t}$, OMWU incorporates a predictive “hint” about the next payoff. Specifically, it forms the optimistic estimate $m_{t}\;=\;2\hat{v}_{t}-\hat{v}_{t-1},\text{with}\;\hat{v}_{0}=\mathbf{0}.$ The meta-strategy then updates according to
@@
followed by normalization to ensure $\sigma_{t+1}\in\Delta_{k_{t}-1}$. [^5] In the context of the Optimistic Multiplicative Weights Update (OMWU) algorithm, $\eta_{t}$ is the step size, often referred to as the learning rate or gain parameter. This optimistic step yields stronger theoretical guarantees: regret now scales with the cumulative variation of the payoff vectors rather than the iteration horizon $T$.
@@ 3.4 Finding New Strategies with a Bandit Oracle
Solving the restricted meta-game alone is insufficient; progress requires introducing new, challenging policies into the population. Gems casts this search as a multi-armed bandit problem (Robbins (1952)). The “arms” comprise a finite pool of candidate latent codes $\Lambda_{t}$, and pulling an arm corresponds to evaluating the policy $G_{\theta}(z)$ against the current meta-strategy $\sigma_{t}$.
@@
Gems estimates this value, $\hat{\mu}_{t}(z)$, together with its empirical variance $\widehat{\mathrm{Var}}_{t}(z)$ via Monte Carlo rollouts. To balance exploration and exploitation efficiently, it employs an empirical-Bernstein Upper Confidence Bound (EB-UCB) (Maurer & Pontil (2009)), which leverages variance information for tighter bounds. A Jacobian penalty encourages smoothness in the generator’s latent space, aiding optimization, where $JG_{\theta}(z)$ denotes the Jacobian matrix of the generator with respect to its input $z$. The score assigned to candidate $z$ is
$\mathrm{UCB}^{\mathrm{EB}}_{t}(z)\;=\;\hat{\mu}_{t}(z)+\sqrt{\frac{2\,\widehat{\mathrm{Var}}_{t}(z)\,\ln(3/\delta_{t})}{n_{z}m}}+\frac{3\ln(3/\delta_{t})}{n_{z}m-1}\;-\;\lambda_{J}\|JG_{\theta}(z)\|_{F}^{2},$ (7)
where $\delta_{t}=t^{-2}$ is a decaying confidence parameter and $\lambda_{t}$ $\geq$ 0 is a penalty coefficient. Gems selects:
$z_{t}^{\star}\;=\;\arg\max_{z\in\Lambda_{t}}\mathrm{UCB}^{\mathrm{EB}}_{t}(z),$ (8)
and augments the anchor set:
$Z_{t+1}\leftarrow Z_{t}\cup\{z_{t}^{\star}\}.$ (9)